OpenStudy (anonymous):

-x + 3y = 9 how could you turn this into a slope intercept form equation. PLZ show step-by-step

OpenStudy (lgbasallote):

first step..put y on the left side ONLY \[3y = x + 9\] next step.. divide everything by the coefficient of y (which in this case is 3) \[\frac{\cancel{3} y}{\cancel{3}} = \frac{x+9}{3}\]

OpenStudy (lgbasallote):

you can also rewrite this as \[ y = \frac{1}{3}x + 3\]

OpenStudy (anonymous):

-x+3y=9 +x +x ____________ 3y = x+9 divide 3 on both side you get y= 1/3x+3<------ answer